OpenStudy (anonymous):

is this --> f(x)=8x-4x^2 graph wider, narrower, or the same as the y=x^2 graphed??? help me.

jimthompson5910 (jim_thompson5910):

f(x)=8x-4x^2 f(x)=-4x^2+8x Since the leading coefficient is -4, which is greater than 1 (the x^2 coefficient of y=x^2), this means that the graph is mirrored over the x axis and it is vertically stretched. This makes the graph more narrow So f(x)=8x-4x^2 is more narrower than y=x^2

OpenStudy (anonymous):

Obviously it's not the same width because those are not the same functions! Graphing them by hand, it seems 8x-4x^2 is narrower. 2nd opinion?

jimthompson5910 (jim_thompson5910):

oops I meant to say that -4 has a larger absolute value than 1

OpenStudy (anonymous):

thanks thanks thanks, i suck at graphing. :)

OpenStudy (anonymous):

There you go.